# What One Cannot, Two Can: Two-Layer Transformers Provably Represent Induction Heads on Any-Order Markov Chains

### v1

@@ Abstract
In-context learning (ICL) is a hallmark capability of transformers, through which trained models learn to adapt to new tasks by leveraging information from the input context. Prior work has shown that ICL emerges in transformers due to the presence of special circuits called induction heads. Given the equivalence between induction heads and conditional $k$-grams, a recent line of work modeling sequential inputs as Markov processes has revealed the fundamental impact of model depth on its ICL capabilities: while a two-layer transformer can efficiently represent a conditional $1$-gram model, its single-layer counterpart cannot solve the task unless it is exponentially large. However, for higher order Markov sources, the best known constructions require at least three layers (each with a single attention head) — leaving open the question: can a two-layer single-head transformer represent any $k^{\text{th}}$-order Markov process? In this paper, we precisely address this and theoretically show that a two-layer transformer with one head per layer can indeed represent any conditional $k$-gram. Thus, our result provides the tightest known characterization of the interplay between transformer depth and Markov order for ICL. Building on this, we further analyze the learning dynamics of our two-layer construction, focusing on a simplified variant for first-order Markov chains, illustrating how effective in-context representations emerge during training. Together, these results deepen our current understanding of transformer-based ICL and illustrate how even shallow architectures can surprisingly exhibit strong ICL capabilities on structured sequence modeling tasks. Code is available at https://anonymous.4open.science/r/markov-llm-depth-icl-63F0.
@@ 1 Introduction
Transformers, powered by the attention mechanism, have emerged as the dominant architecture in machine learning, achieving state-of-the-art performance across a wide range of domains, including natural language processing Brown et al. (2020), computer vision Vaswani et al. (2017), and complex reasoning tasks OpenAI (2023); Yang et al. (2024). A key factor underpinning this success is their ability to efficiently model sequences and perform in-context learning (ICL)—adapting to unseen tasks during inference by leveraging the relevant input context Olsson et al. (2022). It is well known that ICL emerges in transformers due to the presence of special circuits called induction heads Olsson et al. (2022); Singh et al. (2024).
Intuitively, these circuits enable transformers to implement a “copy-and-match” mechanism by copying earlier tokens from the input and matching them with the desired context for next-token prediction. For example, a first-order induction head mimics the functionality $\ldots,A,B,\ldots,A]\to B$. Capitalizing on the connection between induction heads and conditional $k$-gram models, a recent active line of work has leveraged $k^{\text{th}}$-order Markov chains as a simple yet powerful framework to analyze how transformers learn induction heads, with the framework referred to as Markov-ICL Edelman et al. (2024); Nichani et al. (2024); Rajaraman et al. (2024); Chen et al. (2024); Bietti et al. (2023). In particular, using Markovian input sequences, the goal is to understand how transformers learn to represent conditional $k$-grams–equivalent to $k^{\text{th}}$-order induction heads (Def. 2.1).
Building upon the Markov-ICL setting, Bietti et al. (2023); Edelman et al. (2024), and Nichani et al. (2024) demonstrate that a conditional $1$-gram model can be efficiently represented by a two-layer (i.e., two attention layers), single-head transformer. In contrast, Sanford et al. (2024) show that a one-layer, single-head counterpart cannot solve this task unless its hidden dimension is exponentially larger than that of the two-layer model. On the other hand, for higher-order processes, Edelman et al. (2024) and Nichani et al. (2024) argue that low-depth transformers need the number of heads scaling linearly in $k$, in order to learn $k^{\text{th}}$-order Markov processes. However, Rajaraman et al. (2024) establish a surprising result that a three-layer, single-head transformer can represent the conditional $k$-gram model for any $k\geq 1$, and thereby learn $k^{\text{th}}$-order Markov models in-context.
@@
Bietti et al. (2023) 11 22-layer, 11-head $9d^{2}+dS+Td$
Edelman et al. (2024) 11 22-layer, 11-head $21S^{2}+3TS$
Nichani et al. (2024) $k$ 22-layer, $k$-head $(k+(k+1)^{2})(S+T)^{2}+T(S+T)$
Rajaraman et al. (2024) $k$ 33-layer, 11-head $15(3+6S)^{2}+(S+T)(3+6S)$
Ours $k$ 22-layer, 11-head $9(3+6S)^{2}+(S+T)(3+6S)$
Table 1: Comparison with prior works on transformers with Markov-ICL. Here $k\geq 1$ is the Markov order, $T$ the sequence length, and $S$ the state-space (vocabulary) size. Note that Bietti et al. (2023) do not explicitly state what $d$ is and assume it to be large enough. To the best of our knowledge, our work obtains the tightest known characterization of transformer depth and Markov order for higher-order processes. The bit precision for our representation results is the same as that of Rajaraman et al. (2024): it suffices to have $\Omega(\log(T)+k)$ bits per parameter, with $\mathcal{O}(1/T)$ additive approximation error.
Comparison with other Markov-ICL works vis-a-vis transformer architecture. In this paper, as we focus on higher-order processes, we consider the general transformer architecture (Sec. 2.2) closer to real-world models with relative positional encodings Shaw et al. (2018), softmax attention Vaswani et al. (2017), MLPs, and layer normalization Ba et al. (2016). For first-order processes, past literature has considered simplified variants: Bietti et al. (2023) use frozen and fixed random embeddings with no MLPs in the first layer, Edelman et al. (2024) consider attention-only models with no MLPs and embeddings, and Nichani et al. (2024) study a variant called disentangled transformer, splitting the residual and attention streams, with the hidden dimension scaling with depth. While our gradient descent result for first-order processes (Sec. 5) is similar to that of Nichani et al. (2024), the architecture is fundamentally different.
@@
Scalars are denoted by lowercase italic letters (e.g., $x,y$), vectors by lowercase bold letters (e.g., $\mathbf{x},\mathbf{y}$), and matrices by uppercase bold letters (e.g., $\mathbf{A},\mathbf{B}$). The matrices $\mathbf{0}_{m\times n}$ and $\mathbf{1}_{m\times n}$ represent the all-zeros and all-ones matrices in $\mathbb{R}^{m\times n}$, respectively. The indicator function is denoted either by $\mathbb{I}(\cdot)$ or $\mathbbm{1}_{(\cdot)}$, depending on the context. The vocabulary $\mathcal{S}=\{1,2,\ldots,S\}$ of cardinality $|\mathcal{S}|=S$ denotes the finite state-space of the Markov Chains studied in this paper. A sequence $x_{0:T}$ is defined as $x_{0:T}\coloneqq x_{0},x_{1},\ldots,x_{T}$, where $T\in\mathbb{N}$. For any $x\in\mathcal{S}$, its one-hot embedding is denoted by $e_{x}^{S}\in\mathbb{R}^{S}$. $\mathcal{O}(\cdot)$ is the big O notation.
@@ 1.1 Related Work
In recent years, transformers have been widely studied from both theoretical and mechanistic perspectives Weiss et al. (2021); Li et al. (2023); Nguyen (2024). Foundational results Yun et al. (2020); Pérez et al. (2021); Wei et al. (2022a) have established their universality and Turing-completeness, with further work examining their ability to model formal languages Bhattamishra et al. (2020); Liu et al. (2023) and implement algorithmic behaviors such as induction heads Olsson et al. (2022). Concurrently, recent studies have explored how transformers perform ICL, including their learning dynamics and phase transitions Singh et al. (2024); Wei et al. (2022b); Bai et al. (2023); Collins et al. (2024); Ren and Liu (2024); Bai et al. (2023); Akyürek et al. (2023); Hoogland et al. (2024); Lin and Lee (2024). On this front, our work is most closely related to recent works studying Markov-ICL, where the sequential inputs are modeled as Markov processes. In Bietti et al. (2023) and Edelman et al. (2024), the authors discover a stage-wise learning procedure for first-order sources, and show that a $2$-layer, 11-head transformer can represent a first-order induction head. Makkuva et al. (2024a, b) study the optimization landscape and learning dynamics of a $1$-layer, 11-head model for a first-order global Markov chain, unveiling the significance of weight-tying and Markov switching probabilities. Nichani et al. (2024) studies the learning dynamics of $2$-layer, 11-head disentangled transformer, illustrating it learns to implement a first-order induction head. For higher-order sources, Nichani et al. (2024) constructs a $2$-layer, $k$-head architecture, whereas Rajaraman et al. (2024) shows that a $3$-layer, 11-head model suffices to represent a conditional $k$-gram model. In contrast, we show that a $2$-layer, 11-head transformer suffices to represent arbitrary-order Markov models in-context.
@@ 2.1 Markov Processes and the Conditional $k$-gram Model
(a) The conditional $k$-gram model (Def. 2.1). It first (1) identifies the positions in the sequence where the preceding $k$ tokens match the current context (blue), and (2) returns the empirical distribution over the symbols (green) that follow these matched positions.
(b) $k\textsuperscript{th}$-order induction head (Def. 3) The attention pattern concentrates on positions in the sequence where the preceding $k$ tokens exactly match the final $k$ tokens. These are the positions the model attends to most strongly, as they reflect the same context as the current prediction target.
A $k\textsuperscript{th}$-order Markov process is a time-homogeneous stochastic process over a finite state space, where the probability of the next symbol depends only on the most recent $k$ symbols in the sequence Norris (1998). More formally, it is defined as follows:
@@
Definition 2. (Conditional k{k}-gram model) Given a sequence $x_{0:T}\coloneqq x_{0},x_{1},\cdots,x_{T}$ in $\mathcal{S}^{T+1}$, the conditional $k$-gram model $\widehat{\pi}_{k}(s\mid x_{T},x_{T-1},\ldots,x_{0})$ computes the empirical probability of observing a symbol $s\in\mathcal{S}$ matching the last $k$ tokens, i.e. $\widehat{\pi}_{k}(s\mid x_{0},x_{1},\cdots,x_{T})\coloneqq\frac{\sum\limits_{i=k}^{T}\mathbb{I}(x_{i}=s,\,x_{i-1}=x_{T},\,x_{i-2}=x_{T-1},\ldots,\,x_{i-k}=x_{T-k+1})}{\sum\limits_{i=k}^{T}\mathbb{I}(x_{i-1}=x_{T},\,x_{i-2}=x_{T-1},\ldots,\,x_{i-k}=x_{T-k+1})},$
@@ 2.2 Transformer Model
While there exist several attention-based transformer architectures in literature, we adopt the formulation described in Shaw et al. (2018) and Dai et al. (2019), which employs relative positional encodings. Each layer in our transformer comprises softmax-based self-attention, optionally followed by a multilayer perceptron (MLP). Following the construction outlined in Rajaraman et al. (2024), we define a $L$-layer transformer with multi-headed attention as follows:
Definition 3. (Multi-head Attention Transformer) Input $x_{0:T}$, number of layers $L$, number of attention heads per layer $H_{\ell}$ Output distribution $P_{\theta}(\cdot\mid x_{0:T})$ $\textbf{x}_{n}^{(1)}\leftarrow\texttt{Emb}(x_{n})$ for $n=0,1,\dots,T$ for $\ell=1$ to $L$ do   for $n=0$ to $T$ do    for $h=1$ to $H_{\ell}$ do      $\widetilde{\textbf{x}}_{n}^{(\ell,\texttt{head:}h)}\coloneqq\sum_{i=0}^{n}\operatorname{att}_{n,i}^{(\ell,\texttt{head:}h)}\cdot\left(\mathbf{W}_{V}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}+\mathbf{p}_{n-i}^{(\ell,\texttt{head:}h,V)}\right)$    end for    $\widetilde{\textbf{x}}_{n}^{(\ell+1)}\coloneqq\textbf{x}_{n}^{(\ell)}+\sum_{h=1}^{H_{\ell}}\widetilde{\textbf{x}}_{n}^{(\ell,\texttt{head:}h)}$ $\triangleright$ Residual connection    $\textbf{x}_{n}^{(\ell+1)}\coloneqq\texttt{MLP}(\textbf{x}_{n}^{(\ell)},\widetilde{\textbf{x}}_{n}^{(\ell+1)})$ $\triangleright$ MLP with layer normalization and skip connections   end for end for $\operatorname{logits}_{T}\coloneqq\mathbf{W}_{o}\mathbf{x}_{T}^{(L+1)}+b_{o}$ $P_{\theta}(\cdot\mid x_{0:T})\coloneqq f(\operatorname{logits}_{T})$,
where the attention weights
$\operatorname{att}_{n,i}^{(\ell,\texttt{head:}h)}\coloneqq\texttt{softmax}_{i}\left(\left\langle\mathbf{W}_{K}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}+\mathbf{p}_{n-i}^{(\ell,\texttt{head:}h,K)},\;\mathbf{W}_{Q}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}\right\rangle\right).$

### v2

@@ Abstract
In-context learning (ICL) is a hallmark capability of transformers, through which trained models learn to adapt to new tasks by leveraging information from the input context. Prior work has shown that ICL emerges in transformers due to the presence of special circuits called induction heads. Given the equivalence between induction heads and conditional $k$-grams, a recent line of work modeling sequential inputs as Markov processes has revealed the fundamental impact of model depth on its ICL capabilities: while a two-layer transformer can efficiently represent a conditional $1$-gram model, its single-layer counterpart cannot solve the task unless it is exponentially large. However, for higher order Markov sources, the best known constructions require at least three layers (each with a single attention head) — leaving open the question: can a two-layer single-head transformer represent any $k^{\text{th}}$-order Markov process? In this paper, we precisely address this and theoretically show that a two-layer transformer with one head per layer can indeed represent any conditional $k$-gram. Thus, our result provides the tightest known characterization of the interplay between transformer depth and Markov order for ICL. Building on this, we further analyze the learning dynamics of our two-layer construction, focusing on a simplified variant for first-order Markov chains, illustrating how effective in-context representations emerge during training. Together, these results deepen our current understanding of transformer-based ICL and illustrate how even shallow architectures can surprisingly exhibit strong ICL capabilities on structured sequence modeling tasks. Code is available at the link.
@@ 1 Introduction
“A complex system that works is invariably found to have evolved from a simple system that worked.” John Gall, Systemantics (1975)
Transformers, powered by the attention mechanism, have emerged as the dominant architecture in machine learning, achieving state-of-the-art performance across a wide range of domains, including natural language processing [brown2020language], computer vision [vaswani2017attention], and complex reasoning tasks [openai2023gpt4, yang2024qwen2]. A key factor underpinning this success is their ability to efficiently model sequences and perform in-context learning (ICL)—adapting to unseen tasks during inference by leveraging the relevant input context [olsson2022context]. It is well known that ICL emerges in transformers due to the presence of special circuits called induction heads [olsson2022context, singh2024needs].
Intuitively, these circuits enable transformers to implement a “copy-and-match” mechanism by copying earlier tokens from the input and matching them with the desired context for next-token prediction. For example, a first-order induction head mimics the functionality $\ldots,A,B,\ldots,A]\to B$. Capitalizing on the connection between induction heads and conditional $k$-gram models, a recent active line of work has leveraged $k^{\text{th}}$-order Markov chains as a simple yet powerful framework to analyze how transformers learn induction heads, with the framework referred to as Markov-ICL [edelman2024evolution, nichani2024how, rajaraman2024transformers, chen2024unveiling, bietti2023birth]. In particular, using Markovian input sequences, the goal is to understand how transformers learn to represent conditional $k$-grams–equivalent to $k^{\text{th}}$-order induction heads (Subsec. 2.1).
Building upon the Markov-ICL setting, bietti2023birth, edelman2024evolution, and nichani2024how demonstrate that a conditional $1$-gram model can be efficiently represented by a two-layer (i.e., two attention layers), single-head transformer. In contrast, sanford2024one show that a one-layer, single-head counterpart cannot solve this task unless its hidden dimension is exponentially larger than that of the two-layer model. On the other hand, for higher-order processes, edelman2024evolution and nichani2024how argue that low-depth transformers need the number of heads scaling linearly in $k$, in order to learn $k^{\text{th}}$-order Markov processes. However, rajaraman2024transformers establish a surprising result that a three-layer, single-head transformer can represent the conditional $k$-gram model for any $k\geq 1$, and thereby learn $k^{\text{th}}$-order Markov models in-context.
@@
Comparison with other Markov-ICL works vis-a-vis transformer architecture.
In this paper, as we focus on higher-order processes, we consider the general transformer architecture (Subsec. 2.2) closer to real-world models with relative positional encodings [shaw2018self], softmax attention [vaswani2017attention], MLPs, and layer normalization [ba2016layernormalization]. For first-order processes, past literature has considered simplified variants: bietti2023birth use frozen and fixed random embeddings with no MLPs in the first layer, edelman2024evolution consider attention-only models with no MLPs and embeddings, and nichani2024how study a variant called disentangled transformer, splitting the residual and attention streams, with the hidden dimension scaling with depth. While our gradient descent result for first-order processes (Sec. 5) is similar to that of nichani2024how, the architecture is fundamentally different. In Tab. 1, we compare the parameter counts of our construction against alternative constructions. Our results show that, as both the order and sequence length increase, our architecture remains the most compact in terms of parameter efficiency.
@@
bietti2023birth 11 22-layer, 11-head $9d^{2}+dS+Td$
edelman2024evolution 11 22-layer, 11-head $21S^{2}+3TS$
nichani2024how $k$ 22-layer, $k$-head $(k+(k+1)^{2})(S+T)^{2}+T(S+T)$
rajaraman2024transformers $k$ 33-layer, 11-head $15(6S+3)^{2}+(6S+3)(3T+2S+10)$
Ours $k$ 22-layer, 11-head $9(6S+3)^{2}+(6S+3)(2T+2S+9)$
Table 1: Comparison with prior works on transformers with Markov-ICL. Here $k\geq 1$ is the Markov order, $T$ the sequence length, and $S$ the state-space (vocabulary) size. Note that bietti2023birth do not explicitly state what $d$ is and assume it to be large enough. To the best of our knowledge, our work obtains the tightest known characterization of transformer depth and Markov order for higher-order processes. The bit precision for our representation results is the same as that of rajaraman2024transformers: it suffices to have $\Omega(\log(T)+k)$ bits per parameter, with $\mathcal{O}(1/T)$ additive approximation error. A detailed breakdown of the parameter count computations can be found in App. I. Additional experiments analyzing the effect of varying bit precisions are presented in Subsec. H.4.
@@
Scalars are denoted by lowercase italic letters (e.g., $x,y$), vectors by lowercase bold letters (e.g., $\mathbf{x},\mathbf{y}$), and matrices by uppercase bold letters (e.g., $\mathbf{A},\mathbf{B}$). The matrices $\mathbf{0}_{m\times n}$ and $\mathbf{1}_{m\times n}$ represent the all-zeros and all-ones matrices in $\mathbb{R}^{m\times n}$, respectively. For norms and inner products we use standard notation, i.e., norms are written as $||\cdot||$, with $\mathbf{x}||_{2}$ denoting the Euclidean norm. The inner product in Euclidean space between two vectors $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\langle\mathbf{x},\mathbf{y}\rangle$. The indicator function is denoted either by $\mathbb{I}(\cdot)$ or $\mathbbm{1}_{(\cdot)}$, depending on the context. The vocabulary $\mathcal{S}=\{1,2,\ldots,S\}$ of cardinality $|\mathcal{S}|=S$ denotes the finite state-space of the Markov Chains studied in this paper. A sequence $x_{0:T}$ is defined as $x_{0:T}\coloneqq x_{0},x_{1},\ldots,x_{T}$, where $T\in\mathbb{N}$. For any $x\in\mathcal{S}$, its one-hot embedding is denoted by $e_{x}^{S}\in\mathbb{R}^{S}$. $\mathcal{O}(\cdot)$ denotes the big O notation.
@@ 1.1 Related Work
In recent years, transformers have been widely studied from both theoretical and mechanistic perspectives [weiss2021thinking, li2023transformers, nguyen2024understanding]. Foundational results [Yun2020seq2seq, perez2021turing, wei2022turing-approx] have established their universality and Turing-completeness, with further work examining their ability to model formal languages [bhattamishra2020ability, liu2023transformers] and implement algorithmic behaviors such as induction heads [olsson2022context]. Concurrently, recent studies have explored how transformers perform ICL, including their learning dynamics and phase transitions [singh2024needs, wei2022chain, bai2023transformers, collins2024context, ren2024towards, bai2023transformers, akyürek2023learning, hoogland2024developmental, lin2024dual]. On this front, our work is most closely related to recent works studying Markov-ICL, where the sequential inputs are modeled as Markov processes. In bietti2023birth and edelman2024evolution, the authors discover a stage-wise learning procedure for first-order sources, and show that a $2$-layer, 11-head transformer can represent a first-order induction head. makkuva2024attention, makkuva2024local study the optimization landscape and learning dynamics of a $1$-layer, 11-head model for a first-order global Markov chain, unveiling the significance of weight-tying and Markov switching probabilities. nichani2024how studies the learning dynamics of $2$-layer, 11-head disentangled transformer, illustrating it learns to implement a first-order induction head. For higher-order sources, nichani2024how constructs a $2$-layer, $k$-head architecture, whereas rajaraman2024transformers shows that a $3$-layer, 11-head model suffices to represent a conditional $k$-gram model. In contrast, we show that a $2$-layer, 11-head transformer suffices to represent arbitrary-order Markov models in-context.
@@ 2.1 Markov Processes and the Conditional $k$-gram Model
(a) The conditional $k$-gram model (Subsec. 2.1). It first (1) identifies the positions in the sequence where the preceding $k$ tokens match the current context (blue), and (2) returns the empirical distribution over the symbols (green) that follow these matched positions.
(b) $k\textsuperscript{th}$-order induction head (Sec. 3) The attention pattern concentrates on positions in the sequence where the preceding $k$ tokens exactly match the final $k$ tokens. These are the positions the model attends to most strongly, as they reflect the same context as the current prediction target.
A $k\textsuperscript{th}$-order Markov process is a time-homogeneous stochastic process over a finite state space, where the probability of the next symbol depends only on the most recent $k$ symbols in the sequence [norris1998markov]. More formally, it is defined as follows:
@@
Definition 2. (Conditional k{k}-gram model) Given a sequence $x_{0:T}\coloneqq x_{0},x_{1},\cdots,x_{T}$ in $\mathcal{S}^{T+1}$, the conditional $k$-gram model $\widehat{\pi}_{k}(s\mid x_{T},x_{T-1},\ldots,x_{0})$ computes the empirical probability of observing a symbol $s\in\mathcal{S}$ matching the last $k$ tokens, i.e. $\widehat{\pi}_{k}(s\mid x_{0},x_{1},\cdots,x_{T})\coloneqq\frac{\sum\limits_{i=k}^{T}\mathbb{I}(x_{i}=s,\,x_{i-1}=x_{T},\,x_{i-2}=x_{T-1},\ldots,\,x_{i-k}=x_{T-k+1})}{\sum\limits_{i=k}^{T}\mathbb{I}(x_{i-1}=x_{T},\,x_{i-2}=x_{T-1},\ldots,\,x_{i-k}=x_{T-k+1})}$
@@ 2.2 Transformer Model
While there exist several attention-based transformer architectures in literature, we adopt the formulation described in shaw2018self and dai2019transformerxlattentivelanguagemodels, which employs relative positional encodings. Each layer in our transformer comprises softmax-based self-attention, optionally followed by a multilayer perceptron (MLP). Following the construction outlined in rajaraman2024transformers, we define a $L$-layer transformer with multi-headed attention as follows:
Definition 3. (Multi-head Attention Transformer) Input $x_{0:T}$, number of layers $L$, number of attention heads per layer $H_{\ell}$ Output distribution $P_{\theta}(\cdot\mid x_{0:T})$ $\textbf{x}_{n}^{(1)}\leftarrow\texttt{Emb}(x_{n})$ for $n=0,1,\dots,T$ for $\ell=1$ to $L$ do   for $n=0$ to $T$ do    for $h=1$ to $H_{\ell}$ do      $\widetilde{\textbf{x}}_{n}^{(\ell,\texttt{head:}h)}\coloneqq\sum_{i=0}^{n}\operatorname{att}_{n,i}^{(\ell,\texttt{head:}h)}\cdot\left(\mathbf{W}_{V}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}+\mathbf{p}_{n-i}^{(\ell,\texttt{head:}h,V)}\right)$    end for    $\widetilde{\textbf{x}}_{n}^{(\ell+1)}\coloneqq\textbf{x}_{n}^{(\ell)}+\sum_{h=1}^{H_{\ell}}\widetilde{\textbf{x}}_{n}^{(\ell,\texttt{head:}h)}$ $\triangleright$ Residual connection    $\textbf{x}_{n}^{(\ell+1)}\coloneqq\texttt{MLP}(\textbf{x}_{n}^{(\ell)},\widetilde{\textbf{x}}_{n}^{(\ell+1)})$ $\triangleright$ MLP with layer normalization and skip connections   end for end for $\operatorname{logits}_{T}\coloneqq\mathbf{W}_{o}\mathbf{x}_{T}^{(L+1)}+b_{o}$ $P_{\theta}(\cdot\mid x_{0:T})\coloneqq f(\operatorname{logits}_{T})$, Where the attention weights are defined as: $\operatorname{att}_{n,i}^{(\ell,\texttt{head:}h)}\coloneqq\texttt{softmax}_{i}\left(\left\langle\mathbf{W}_{K}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}+\mathbf{p}_{n-i}^{(\ell,\texttt{head:}h,K)},\;\mathbf{W}_{Q}^{(\ell,\texttt{head:}h)}\textbf{x}_{i}^{(\ell)}\right\rangle\right).$